# Comprehensive Evaluation of Prototype Neural Networks

### v1

@@ 5.1 Output-Completeness Evaluation
The output-completeness evaluation focuses on the visualization method’s ability to highlight all relevant parts in the input image. To evaluate the output completeness of the PRP method, we consider the study by Sacha et al. [57], which evaluates the spatial misalignment of visualization methods. The introduced metrics measure the location change in the visualization VLC, the prototype activation change PSC, and the prototype rank change PRC when a perturbation is applied to the input image. We extend this list by including the activation change of the visualization VAC, the location change of the maximum activation in the feature map PLC, as well as the general location and activation change of the similarity pattern, PALC and PAC, respectively. Some metrics from Sacha et al. [57] have been renamed for consistency in terminology.
Consider in image and the perturbed counterpart , created based on the visualization of the prototype . Let be the saliency map of prototype . Each saliency map is pre-processed and only contains pixels with a relevance value above the 95th percentile. Let be the bounding box around this activated region of the saliency map. Further, let be the similarity map and the similarity score of the prototype. We perturb an image to by adding a Gaussian noise mask with a standard deviation of to the pixels around the bounding box.
If the visualization method correctly highlights the relevant parts of the prototype, then we expect minimal changes in the model, resulting in low scores for these metrics.
@@
The final metric from Sacha et al. [57] is the prototype rank change PRC. This metric measures the prototype’s position change in the similarity score ranking of all prototypes.
@@ 5.2 Continuity Evaluation
The continuity evaluation assesses the stability of prototypes under different augmentations. We focus on the study by Hoffmann et al. [48], which assesses prototype stability under noise and compression artifacts, and the studies by Rymarczyk et al. [33] and Nauta et al. [53], which examine the influence of photometric augmentations. We train our model using geometric and photometric augmentations to promote continuity during training, as recommended by Nauta et al. [17]. In the evaluation, we perturb the input image to with photometric and noise augmentations. The individual augmentations are applied all at once with the following settings: brightness , contrast , saturation , hue , noise , JPEG quality , and blur using a kernel.
We utilize the metrics from the output-completeness evaluation to evaluate the changes in location and activation within the model. The VLC and VAC scores are excluded, as they focus on assessing the visualization method instead of a prototype. We also introduce the following new metrics. The classification activation change CAC measures the change in prediction. Let be the number of classes and be the output vector of the model for an image .
@@ 5.3 Contrastivity
We use the average inter-class distance metrics from Wang et al. [58] and also employ the average intra-class distance case and respectively. Following the recommendations from Nauta et al. [17], we use the PLC and PALC metrics to assess the contrastivity of prototype locations in the feature map and introduce a new metric to evaluate the activation discriminativeness of prototypes.
The original study from Wang et al. [58] does not state the specific formula of the metrics, so the following calculations are our interpretation of the metrics. The average inter-class prototype distance from Wang et al. [58] is the average cosine distance between prototypes of different classes and is calculated as follows:
We calculate the average inter-class feature distance from Wang et al. [58] in the same way using the feature vector set instead of the prototype set . To evaluate the contrastivity between prototypes from the same class, we introduce the average intra-class prototype distance . The metric measures the average cosine distance of a prototype to all other prototypes of the same class and is calculated as follows:
We also extend this metric to the feature vector case introducing the average intra-class feature distance , by switching the prototype set with the feature vector set . To measure the class discriminativeness of a prototype, we compute the Shannon entropy of the similarity scores. Let be the set of similarity scores of a prototype over all test images in . We first use a max normalization to normalize the similarity scores and then calculate the histogram of the similarity scores with bins. The discriminativeness of the prototype is then calculated as follows:
@@ 5.4 Covariate Complexity
To assess the overlap between prototypes and objects, Wang et al. [58] introduces the content heatmap metric. We renamed this metric to to be consistent with our naming scheme. The metric calculates the intersection of the prototype visualization and the object mask. Let be the object mask of image . The metric is calculated as follows:
@@
To measure the alignment of prototypes with specific object parts, we adapted the consistency score from Huang et al. [49]. The coverage of an object part is calculated by computing a histogram of the object parts covered by a prototype over the test set . This is done by computing the bounding box and adding all object parts within the bounding box to the histogram of that prototype. After processing every image in the test set, we normalize the histograms. This results in a vector with object part percentages, denoted as . The consistency of the prototype is the average coverage measure over the vector and is calculated as follows:
In addition, we introduce the metric as a counterpart to the . This metric assesses the area of the activated region in the saliency map that does not overlap with the object. The metric is calculated as follows:
@@ 5.5 Compactness
Then, we count the number of prototypes with a normalized similarity score above a threshold , which is set to :
@@ 7 Results
The general performance, compactness, contrastivity, and continuity evaluations are conducted on all datasets. The output-completeness and complexity evaluations are conducted exclusively on the CUB200 dataset. The output-completeness evaluation exclusively assesses the employed visualization method, in our case the computationally intensive PRP method; thus, this method was evaluated only on the CUB200 dataset. The complexity evaluation is limited to the CUB200 dataset due to the absence of object masks and part annotations in the other datasets.
@@ 7.1 General and Compactness
Table 2 presents the outcomes of our evaluations on general performance and compactness. The results obtained from ProtoPNet indicate that pruning adversely affects performance. The ProtoPool model exhibits no definitive ranking trend, frequently performing slightly worse than the ProtoPNet models but demonstrating an advantage on the Cars196 and AWA2 datasets. The PIPNet model exhibits the best performance across all datasets.
Notably, all networks exhibit a considerable performance improvement when comparing overall accuracy to top-3 accuracy, showing potential gains if the models are further tuned, particularly on the fine-grained datasets CUB200 and Cars196, with each containing around 200 classes. The results from the NICO dataset suggest that all networks are similarly influenced by the distribution shift, with no significant performance losses. Furthermore, all models show a significant performance reduction on the multi label dataset AWA2, suggesting that learning attributes like strides from multiple animal classes presents significant challenges. This is followed by the fine-grained datasets, where the CUB200 dataset is more challenging than Cars196. In contrast, the general classification task from the NICO dataset appears to be the least difficult.
| CUB200 | Entropy | ||||||
| ProtoPNet | 1.95 0.28 | 39.83 2.08 | 0.00 0.00 | 0.00 0.00 | 0.01 0.00 | 0.01 0.00 | 0.99 0.00 |
| ProtoPNet P | 1.92 0.28 | 39.05 2.45 | 0.00 0.00 | 0.00 0.00 | 0.01 0.00 | 0.01 0.00 | 0.99 0.00 |
| ProtoPool | 0.71 0.06 | 15.60 2.51 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.91 0.01 |
| PIPNet | 3.99 0.06 | 100.00 0.00 | - | 1.00 0.00 | - | 0.96 0.00 | 0.34 0.01 |
| Cars196 | Entropy | ||||||
| ProtoPNet | 2.02 0.15 | 59.82 3.26 | 0.00 0.00 | 0.00 0.00 | 0.02 0.00 | 0.02 0.00 | 0.96 0.01 |
| ProtoPNet P | 1.94 0.10 | 58.40 2.35 | 0.00 0.00 | 0.00 0.00 | 0.02 0.00 | 0.02 0.00 | 0.96 0.01 |
| ProtoPool | 1.69 0.88 | 40.94 21.70 | 0.01 0.00 | 0.01 0.00 | 0.01 0.01 | 0.01 0.01 | 0.86 0.05 |
| PIPNet | 4.00 0.04 | 100.00 0.00 | - | 1.00 0.00 | - | 0.99 0.00 | 0.35 0.00 |
| NICO | Entropy | ||||||
| ProtoPNet | 1.56 0.27 | 45.71 7.30 | 0.00 0.00 | 0.00 0.00 | 0.03 0.00 | 0.03 0.00 | 0.54 0.04 |
| ProtoPNet P | 1.55 0.26 | 45.64 7.21 | 0.00 0.00 | 0.00 0.00 | 0.03 0.00 | 0.03 0.00 | 0.53 0.04 |
| ProtoPool | 1.46 0.10 | 34.37 4.21 | 0.10 0.02 | 0.03 0.01 | 0.13 0.02 | 0.16 0.01 | 0.75 0.11 |
| PIPNet | 3.86 0.05 | 100.00 0.00 | - | 1.00 0.00 | - | 0.97 0.01 | 0.35 0.00 |
| AwA2 | Entropy | ||||||
| ProtoPNet | 3.86 0.44 | 85.06 2.72 | 0.01 0.00 | 0.01 0.00 | - | - | 0.88 0.03 |
| ProtoPNet P | 3.87 0.40 | 86.10 2.66 | 0.01 0.00 | 0.01 0.00 | - | - | 0.90 0.03 |
| ProtoPool | 2.87 1.02 | 22.96 3.57 | 0.04 0.03 | 0.04 0.03 | - | - | 0.85 0.05 |
| PIPNet | 3.74 0.02 | 100.00 0.00 | - | 1.00 0.00 | - | - | 0.32 0.00 |
The compactness evaluation shows that the global size of PIPNet remains below its allocated number by a large margin across all datasets. ProtoPool could not achieve a reduction of prototypes, but uses the smallest absolute number of prototypes on all datasets, except NICO. The pruning method employed by ProtoPNet consistently reduces the global size of the original model by a small margin, except on the NICO dataset.
The sparsity results reveal that only PIPNet achieves consistent results on all datasets. The results of other models fluctuate, indicating dataset and tasks dependencies.
We see that the ProtoPNet models do not achieve a satisfactory degree of sparsity on the multi-label classification task (AWA2). The sparsity on the general classification task (NICO) is also significantly lower compared to other models. ProtoPool shows only a noticeably lowered sparsity on the NICO dataset. The pruning of the ProtoPNet models reduces sparsity by a small margin. This indicates that the deleted prototypes are associated with fewer classes in the classification layer compared to the retained prototypes, which is an unexpected finding. The pruning strategy calculates prototype-class pairs based on the distance to feature vectors and retains prototypes that have a clear class assignment. Prototypes retained should, therefore, also display a clearer assignment within the classification layer than the deleted ones.
PIPNet consistently achieves a perfect score due to its hard constraint that ensures only positive weights in the classification layer. The pruning of ProtoPNet reduces the local size of the model on the fine-grained classification tasks, indicating that the deleted prototypes were present in a notable portion of the test images. ProtoPool only achieves the best local size results on the NICO dataset despite the small global size, suggesting that the learned prototypes are complex representations without a clear class correspondence in the other datasets. PIPNet consistently achieves one of the best local size results, except for the NICO dataset. The similar results across tasks and datasets suggest that PIPNet’s local size is mostly dependent on the initial number of prototypes.
@@ 7.2 Contrastivity
The results from the contrastivity evaluation are presented in Table 3. Here, we compared the top-5 most activated prototypes of an image. To increase understanding, some metrics have been labelled with an additional subscript to indicate their different meanings across sections.
PIPNet repeatedly achieves the highest score across all datasets. Its score is also consistently the highest. A perfect score for PIPNet indicates that the architecture design employed in combination with the loss functions is very effective in learning focused prototype activations without any overlap with other prototypes. In contrast, the lower scores in other models show their prototype activations are more dispersed. ProtoPool has the lowest score on all datasets, suggesting that its focused similarity approach was unsuccessful in preventing overlapping activation patterns. A low contrast between prototypes is further confirmed by the low scores, showing that the maximum activations are also close to each other.
PIPNet does not have explicit prototype vectors, rendering the prototype distance metrics inapplicable. However, the feature distances of the PIPNet models exhibit nearly perfect results after the softmax layer of the model. In the AWA2 dataset’s multi-label setting, measuring vector distances between labels is also unfeasible due to the 1: relation between sample images and labels. The other models could not learn a reasonable distinction between prototypes on all datasets. This indicates that the contrastivity between classes is learned differently.
Entropy results show that PIPNet exhibits the least fluctuation in similarity scores of individual prototypes on all datasets. This contrasts with our expectation that prototypes shared among classes should have a higher fluctuation. Furthermore, despite the general low performance on the AWA2 dataset, PIPNet retained its low entropy score in comparison to the other models, indicating that the model architecture actively facilitates the learning of distinct prototype activations. The entropy results from the other models are notably higher. Especially, results on the fine-grained and the AWA2 dataset show high entropy values, indicating a correlation with the overall difficulty of the tasks. In general, the architecture of the other models do not actively promote class distinct prototype activations.
@@ 7.3 Continuity
| CUB200 | CAC | CRC | |||||
| ProtoPNet | 1.20 0.14 | 16.93 1.18 | 10.88 3.94 | 34.42 1.80 | 14.09 1.17 | 6.35 0.58 | 1.87 0.25 |
| ProtoPNet P | 1.20 0.16 | 16.98 1.06 | 9.79 3.47 | 34.20 1.87 | 14.16 1.09 | 6.34 0.59 | 1.72 0.17 |
| ProtoPool | 2.64 0.26 | 49.31 9.99 | 2.38 0.17 | 39.52 4.08 | 17.34 0.97 | 6.71 0.17 | 3.70 0.36 |
| PIPNet | 0.05 0.00 | 0.47 0.09 | 0.94 0.04 | 2.89 0.24 | 3.82 0.40 | 6.49 0.17 | 0.34 0.04 |
| Cars196 | CAC | CRC | |||||
| ProtoPNet | 0.79 0.06 | 20.55 1.15 | 9.77 1.32 | 33.54 2.40 | 14.18 0.75 | 5.84 0.38 | 0.58 0.04 |
| ProtoPNet P | 0.78 0.05 | 20.15 0.99 | 8.20 0.77 | 32.82 2.16 | 14.16 0.73 | 5.88 0.37 | 0.62 0.07 |
| ProtoPool | 1.22 0.36 | 34.92 41.02 | 6.44 3.36 | 34.28 15.10 | 17.28 7.19 | 5.11 0.45 | 1.36 0.14 |
| PIPNet | 0.01 0.00 | 0.05 0.01 | 0.77 0.03 | 0.82 0.09 | 0.98 0.09 | 5.68 0.50 | 0.13 0.01 |
| NICO | CAC | CRC | |||||
| ProtoPNet | 1.83 0.15 | 58.56 6.43 | 2.23 0.13 | 54.81 3.41 | 26.15 2.12 | 27.03 1.66 | 0.28 0.04 |
| ProtoPNet P | 1.82 0.15 | 58.90 6.71 | 2.22 0.11 | 54.84 3.37 | 26.21 2.09 | 27.05 1.65 | 0.27 0.03 |
| ProtoPool | 1.43 0.15 | 51.71 5.35 | 1.02 0.21 | 31.37 1.23 | 18.12 2.20 | 16.42 2.70 | 0.36 0.06 |
| PIPNet | 0.02 0.01 | 0.30 0.10 | 1.34 0.12 | 0.81 0.41 | 1.32 0.56 | 11.05 0.50 | 0.08 0.01 |
| AwA2 | CAC | CRC | |||||
| ProtoPNet | 1.48 0.16 | 57.46 2.10 | 5.09 0.72 | 47.01 2.03 | 15.60 0.52 | 8.24 0.54 | - |
| ProtoPNet P | 1.46 0.15 | 56.61 2.61 | 5.23 0.43 | 46.30 2.29 | 15.35 0.70 | 8.68 0.47 | - |
| ProtoPool | 2.43 0.64 | 10.05 3.49 | 1.45 0.10 | 22.35 7.69 | 13.47 2.79 | 4.02 0.30 | - |
| PIPNet | 0.00 0.00 | 0.11 0.01 | 1.34 0.03 | 0.01 0.01 | 0.16 0.02 | 3.79 0.17 | - |
The continuity evaluation measures the change of a single prototype when augmentations are applied on the input image. We will again use the top-5 most activated prototypes per image. To distinguish the continuity results shown in Table 4 from other sections, we have added subscripts to some metrics.
PIPNet consistently achieves the lowest and scores across all datasets by a considerable margin, suggesting that the prototype locations remain robust against applied augmentations. The comparison of and scores among other models shows that ProtoPNet achieves better results on the fine-grained datasets, and ProtoPool shows no distinct performance correlation across tasks.
Furthermore, PIPNet achieves significantly better and scores on all datasets. This indicates that the model design enhances the robustness of extracted features against augmentations compared to other models. We can also see that the pruning of the ProtoPNet models does not adversely impact these metrics. Indicating that the remaining prototypes have similar continuity properties compared to the deleted prototypes.
There is no clear trend when comparing and scores, which may initially seem counterintuitive, as we would expect that smaller changes in the similarity score would also stabilize the ranking of these scores. In contrast, the results indicate a strong model dependency. PIPNet shows near-perfect similarity score and rank stability on all datasets. ProtoPool achieves better rank stability compared to their performance, and ProtoPNet displays the reverse case.
scores show that classification accuracy is influenced by augmentations across all models, with PIPNet demonstrating higher robustness in the classification process. We see similar performance on the fine-grained datasets. However, ProtoPNet shows increased classification activation fluctuation on the other datasets. Again, there is no clear correlation between and similar to the observations between and scores.
Comparing the and scores does not reveal a clear correlation between changes in similarity scores and prediction scores. This indicates a complex relationship between prototypes and classes, which reduces the interpretability of the classification process.
7.4 Output-Completeness and Covariate Complexity
The results of the Output-Completeness and Covariate Complexity evaluation on the CUB200 dataset are presented in Table 5. The Output-Completeness evaluation measures the change of a single prototype, akin to the continuity evaluation. However, the augmentation is derived from the PRP visualization method. Again, we used the top-5 most activated prototypes per image. Subscripts have been added to metrics that were also used in the previous evaluations.
| Completeness | VLC | VAC | |||||
| ProtoPNet | 42.77 1.18 | 9.17 1.17 | 0.24 0.04 | 3.70 0.31 | 0.84 0.06 | 46.49 1.87 | 4.76 0.11 |
| ProtoPNet P | 42.89 1.03 | 9.13 1.19 | 0.25 0.04 | 3.67 0.31 | 0.78 0.05 | 47.56 1.36 | 4.75 0.12 |
| ProtoPool | 44.90 0.83 | 9.38 0.80 | 1.73 0.26 | 43.97 11.10 | 1.32 0.14 | 95.48 1.37 | 10.16 1.39 |
| PIPNet | 43.47 0.45 | 5.75 0.21 | 0.01 0.00 | 0.03 0.01 | 0.16 0.00 | 0.43 0.08 | 0.59 0.06 |
| Complexity | Object Overlap | Background Overlap | IORD | Consistency | - | - | - |
| ProtoPNet | 10.33 1.64 | 24.70 11.79 | 0.03 0.03 | 42.83 3.27 | |||
| ProtoPNet P | 10.49 1.57 | 23.63 11.21 | 0.03 0.03 | 43.32 4.46 | |||
| ProtoPool | 1.98 0.37 | 84.79 2.71 | -0.06 0.01 | 42.17 2.76 | |||
| PIPNet | 5.25 0.03 | 60.37 0.29 | -0.02 0.00 | 48.39 1.04 |
PIPNet consistently achieves the lowest scores across all metrics evaluating the latent space, often by a significant margin, suggesting that prototype activations in the latent space closely correspond to prototype visualizations in the input space. This stands in contrast with the and scores validating the accuracy of the PRP method, in which the PIPNet model exhibits results comparable to those of the other models. This indicates that the PRP method is generally influenced by noise augmentations of irrelevant pixels.
Comparing the original ProtoPNet model to its pruned variant, we can again observe comparable performance. ProtoPool exhibits similar changes in visualization, as shown by the and scores. However, the ProtoPool model shows notable higher changes in the latent space. This is unexpected considering that the PRP method starts at the maximum activation to backpropagate relevance scores. Changes in the maximum location measured with should result in more drastic changes in the input space. This is an indication that the feature vectors in the latent space are not well separated.
An IORD score near zero indicates a similar mean activation inside and outside the object. This suggests that the most important image regions of individual prototypes typically correspond to the object contour lines. This can be seen in Figure 2 in the example images from the CUB200 dataset. However, it is also evident that slightly negative scores are a direct indicator that prototypes focus more on the background, measured by the .
scores suggest that the learned prototypes across all networks predominantly focus on individual parts of the object, as intended. Notably, ProtoPool exhibits the smallest overlap, indicating that its prototypes are the most focused. However, the measured indicates that a significant portion of the prototype activations from ProtoPool is directed towards the background, thereby rendering the assumption of focused prototypes ambiguous. The results from the ProtoPNet models show that the pruning strategy could not lower the metric significantly. This suggests that the learned prototypes are complex mixtures of object and background that cannot easily be split into these two groups.
PIPNet has the highest score, supporting the architectural intention to enhance the semantic alignment of prototypes with object parts relative to prior models. The pruning of ProtoPNet has only slight effects on the , again suggesting that prototypes are a complex mixture of object and background, which makes it difficult to prune a specific subset.
Overall scores remain comparatively low, suggesting that prototypes generally do not align with individual object parts. This significantly affects interpretability, given that the primary objective of these networks is to represent distinct features with prototypes.

### v2

@@ 5.1 Output-Completeness Evaluation
The output-completeness evaluation focuses on the visualization method’s ability to highlight all relevant parts in the input image. To evaluate the output completeness of the PRP method, we consider the study by Sacha et al. [63], which evaluates the spatial misalignment of visualization methods. The introduced metrics measure the location change in the visualization VLC and the prototype activation change PSC when a perturbation is applied to the input image. We extend this list by including the activation change of the visualization VAC, the location change of the maximum activation in the feature map PLC, as well as the general location and activation change of the similarity pattern, PALC and PAC, respectively, thereby achieving comprehensive coverage of the model components (see Figure 1). Some metrics from Sacha et al. [63] have been renamed for consistency in terminology.
Consider an image and the perturbed counterpart , created based on the visualization of the prototype . Let be the saliency map of prototype . Each saliency map is pre-processed and only contains pixels with a relevance value above the 95th percentile. Let be the bounding box around this activated region of the saliency map. We perturb an image to by adding a Gaussian noise mask with a standard deviation of to the pixels around the bounding box; see Figure 5 for a process visualization. Further, let be the similarity map and the similarity score of the prototype.
If the visualization method correctly highlights the relevant parts of the prototype, then we expect minimal changes in the model components (see Figure 1), resulting in low scores for these metrics.
@@ 5.2 Continuity Evaluation
The continuity evaluation assesses the stability of prototypes under different augmentations. We focus on the study by Hoffmann et al. [51], which assesses prototype stability under noise and compression artifacts, and the studies by Rymarczyk et al. [34] and Nauta et al. [57], which examine the influence of photometric augmentations. We train our model using geometric and photometric augmentations to promote continuity during training, as recommended by Nauta et al. [19]. In the evaluation, we perturb the input image to with photometric and noise augmentations. The individual augmentations are applied all at once with the following settings: brightness , contrast , saturation , hue , noise , JPEG quality , and blur using a kernel. A visual presentation of the process is illustrated in Figure 1.
We utilize the metrics from the output-completeness evaluation to evaluate the changes in location and activation within the major components of the models (Figure 1). The VLC and VAC scores are excluded, as they focus on assessing the visualization method instead of a prototype. In addition, we use the prototype rank change PRC from Sacha et al. [63]. This metric measures the prototype’s position change in the similarity score ranking of all prototypes.
We also introduce the following new metrics. The classification activation change CAC measures the change in prediction. Let be the number of classes and be the output vector of the model for an image .
@@ 5.3 Contrastivity
We use the average inter-class distance metrics from Wang et al. [64] and also employ the average intra-class distance case and respectively. Following the recommendations from Nauta et al. [19], we use the PLC and PALC metrics to assess the contrastivity of prototype locations in the feature map and introduce a new metric to evaluate the activation discriminativeness of prototypes, achieving a comprehensive model coverage (see Figure 1).
The original study from Wang et al. [64] does not state the specific formula of the metrics, so the following calculations are our interpretation of the metrics. The average inter-class prototype distance from Wang et al. [64] is the average cosine distance between prototypes of different classes.
We calculate the average inter-class feature distance from Wang et al. [64] in the same way using the feature vector set instead of the prototype set . To evaluate the contrastivity between prototypes from the same class, we introduce the average intra-class prototype distance . The metric measures the average cosine distance of a prototype to all other prototypes of the same class.
We also extend this metric to the feature vector case introducing the average intra-class feature distance , by switching the prototype set with the feature vector set . To measure the class discriminativeness of a prototype, we compute the Shannon entropy of the similarity scores. Let be the set of similarity scores of a prototype over all test images in . We use a max normalization to normalize the similarity scores and then calculate the histogram of the similarity scores with bins.
@@ 5.4 Covariate Complexity
To assess the overlap between prototypes and objects, Wang et al. [64] introduces the content heatmap metric. We renamed this metric to to be consistent with our naming scheme. The metric calculates the intersection of the prototype visualization and the object mask. Let be the object mask of image .
@@
To measure the alignment of prototypes with specific object parts, we adapted the consistency score from Huang et al. [53]. The coverage of an object part is calculated by computing a histogram of the object parts covered by a prototype over the test set . This is done by computing the bounding box and adding all object parts within the bounding box to the histogram of that prototype. After processing every image in the test set, we normalize the histograms. This results in a vector with object part percentages, denoted as . The consistency of the prototype is the average coverage measure over the vector .
In addition, we introduce the metric as a counterpart to the . This metric assesses the area of the activated region in the saliency map that does not overlap with the object.
@@ 5.5 Compactness
Then, we count the number of prototypes with a normalized similarity score above a threshold , which is set to .
@@ 7 Results
The general performance, compactness, contrastivity, and continuity evaluations are conducted on all datasets. The output-completeness and complexity evaluations are conducted exclusively on the CUB200 dataset. The output-completeness evaluation assesses the employed visualization method, in our case the computationally intensive PRP method; thus, this method was evaluated only on the CUB200 dataset. The complexity evaluation is limited to the CUB200 dataset due to the absence of object masks and part annotations in the other datasets.
@@ 7.1 General and Compactness
Table 2 presents the outcomes of our evaluations on general performance and compactness.
The pruned version of ProtoPNet exhibits a performance drop compared to the original model. This indicates that prototypes not clearly associated with a single class, presumably focusing on background regions, still play a significant role in the classification process. In other words, classification in ProtoPNet relies on a complex and sensitive interplay between prototypes, including those that are not class-distinct. Evidence from the compactness analysis supports this interpretation. First, the Local Size shows that a relatively high number of prototypes are active per sample. Second, the NPR score improves after removing background prototypes, suggesting that even though the fine-tuning stage was initialized with zeros (instead of the original –0.5), which should bias towards positive reasoning, the model nonetheless learns intricate negative relations during fine-tuning. Interestingly, sparsity often decreases slightly after pruning, which is counterintuitive. One would expect that the targeted background prototypes would be shared across more classes; thus their removal should increase sparsity. Combined with the minor reduction in global size, these findings suggest that pruning does not simply eliminate “background prototypes” but is strongly influenced by the quality of learned prototypes.
On the CUB200 dataset, ProtoPool struggles to discover meaningful shared prototypes compared to its performance on other datasets, such as Cars196. This is reflected both in performance rankings and compactness measures. Specifically, ProtoPool on CUB200 exhibits a high proportion of negative weights, low sparsity, and a large number of prototypes active per sample, all indicators of a complex classification process. In contrast, when applied to NICO and AWA2, ProtoPool shows improvements in interpretability over ProtoPNet. The Local Size is reduced on NICO, while NPR and Sparsity scores improve on AWA2. These results suggest that shared prototypes may be more advantageous in multi-label settings than in fine-grained single-label classification tasks.
PIPNet achieves the best overall results in both predictive performance and compactness. This illustrates the potential of contrastive learning for prototype-based methods. The outcome is intuitive, as contrastive learning naturally aligns with the goal of learning high-level semantic prototype representation. Furthermore, the hard constraint imposed on the classification layer does not negatively affect performance, indicating that PIPNet does not require a separate fine-tuning phase. In contrast, other models that rely on fine-tuning often learn additional, more complex, relationships to increase predictive performance in this phase. PIPNet is also the only model capable of reducing Global Size during training without the need for external pruning strategies. This highlights the effectiveness of the softmax-based constraint in limiting the number of active prototypes, even while the tanh loss encourages all prototypes to be active at some point during training. Moreover, the consistently small Local Size relative to Global Size indicates that this design robustly limits prototype usage per sample, thereby enhancing interpretability without sacrificing accuracy.
@@ 7.2 Contrastivity
All contrastivity results in Figure 3 were normalized using either fixed metric boundaries or the maximum values observed across models. The metric axes, where lower values indicate better performance, are inverted to consistently increase the coverage area of the more powerful models.
The original and pruned versions of ProtoPNet exhibit highly similar behaviours. Both show weak location contrast in the feature maps and low feature and prototype distance scores, indicating that the learned embedding space is not disentangled. Consequently, the extracted features contain similar information rather than distinct, high-level object-part representations. This lack of disentanglement also reduces class separability in the embedding space, suggesting that the backbone network primarily encodes generic shape and colour information rather than interpretable semantic parts. As a result, the embedding space is comparatively smaller and produces ambiguous prototype activations, as reflected by higher entropy values.
ProtoPool exhibits characteristics similar to ProtoPNet, with an embedding space that appears densely clustered and dominated by generic shape and colour information. This indicates the model’s limiting ability to capture separable high-level features. The attempt to learn shared prototypes seems to exacerbate this issue, compressing the embedding space even further compared to ProtoPNet. However, improvements are observed in the NICO dataset, where inter-class prototype and feature distances are slightly higher. This suggests that in domains where low-level features (e.g., basic shapes) provide stronger class distinctions, such as differentiating between animal species, ProtoPool’s shared prototype approach can be more effective.
PIPNet achieves the highest scores in the feature distance measures. Prototype distances are not taken into account due to the lack of explicit prototype vectors. Unlike the other models, PIPNet does not treat the backbone outputs as embedding vectors; instead, it interprets each feature dimension as a prototype assignment via softmax. Under this formulation, the feature distance can be understood as the degree of overlap between prototype activations. Including the PALC, PLC, and AFD scores, PIPNet shows near-perfect contrast in prototype localization, demonstrating its ability to learn a disentangled feature extraction process. This further indicates that PIPNet successfully learns high-level prototypes representing distinct semantic properties. In other words, the extracted features align with the intended goal of representing meaningful and separable object parts. The Entropy results reinforce this conclusion, showing that PIPNet’s embedding space is more evenly distributed than that of other models. This distribution facilitates discrete prototype-to-class activations, even when prototypes are shared across classes.
@@ 7.3 Continuity
All continuity results in Figure 4 were normalized using either fixed metric boundaries or the maximum values observed across models. Again, we inverted axes of metrics where better performance is indicated by lower scores, so higher performing models cover a larger area in the chart.
The pruned version of ProtoPNet achieves slightly improved PRC scores on the CUB200 and Cars196 datasets compared to the original model. This suggests that pruning successfully removes prototypes more sensitive to low-level perturbations. However, this improvement does not translate into greater robustness at the classification level, as reflected by unchanged CAC and CRC scores. This observation is consistent with the compactness evaluation, which revealed a complex mixture of relations rather than a straightforward classification process. Interestingly, results on the NICO dataset show higher sensitivity to low-level perturbations than on the fine-grained datasets. This indicates that ProtoPNet tends to rely on simpler, low-level features when the overall classification task is less challenging. This underscores the absence of a dedicated mechanism that encourages the learning of high-level features.
The design of ProtoPool also does not include such a mechanism. However, it appears learning shared prototypes promote the extraction of more robust features under the right circumstances. This is reflected in the superior performance on the NICO and AWA2 dataset compared to ProtoPNet. A noteworthy finding is the high PLC score relative to other similarity-map metrics, such as PALC. This suggests that while the location of maximum activation may vary, the broader activation patterns remain relatively stable. Furthermore, ProtoPool performs particularly well on the AWA2 dataset, indicating that the multi-label classification loss contributes to learning more robust shared prototypes. Another key observation is that better PRC scores do not necessarily translate into improved CRC scores. This again highlights the complex and sensitive nature of the classification process from the trained ProtoPool models, which is in agreement with the findings from the compactness evaluation.
PIPNet demonstrates the strongest robustness in both prototype location and activation under low-level image augmentations. This result is expected, as the model employs a contrastive learning strategy that explicitly reduces the influence of such perturbations. Notably, this approach also appears to facilitate the extraction of high-level object-part features, as reflected in the contrastive evaluation. Nevertheless, some fluctuations are observed in classification-focused metrics, particularly in the PRC. The larger variability in PRC compared to the PSC suggests that activations among the top-5 prototypes are highly similar. This is likely a consequence of the softmax layer constraining activations between 0 and 1. The sometimes larger CAC further reveals that the classification layer amplifies these minor differences. Despite some activation fluctuations, the predicted class remains relatively stable, as indicated by the smaller fluctuations in the CRC compared to the CAC.
7.4 Covariate Complexity
All covariate complexity results in Figure 5 were normalized using either fixed metric boundaries or the maximum values observed across models. The metric axes, where lower values indicate higher performance, are inverted to consistently increase the coverage area of better models.
ProtoPNet and its pruned variant produce highly similar results, suggesting that background prototypes were not effectively removed. This finding is consistent with the compactness evaluation, which revealed a complex classification process and indicated that prototypes failed to capture well-defined concepts. Nevertheless, the lower Background Overlap compared to ProtoPool and PIPNet suggests that the class-specific prototypes of the ProtoPNet model focus more on object regions. This is intuitive, as backgrounds often serve as shared concepts across classes, and are therefore well suited to be used as shared prototypes.
The ProtoPool model exhibits the highest Background Overlap, indicating that it is particularly affected by this issue. This observation aligns with its overall performance and compactness results on the CUB200 dataset, which suggest that the learned prototypes lack sufficient class-discriminative information and are of comparatively low quality.
Interestingly, the IORD scores are nearly identical across all models, indicating that, on average, model focus is balanced between object and background regions despite differing levels of Background Overlap. This suggests that prototype activations are relatively uniform, probably due to the 95th percentile crop, leading to similar average scores for both object- and background-overlapping activations. This can also be observed in the example visualization in Figure 6 and 7.
Object-part Consistency is generally low across models, with PIPNet performing slightly better. This indicates that prototypes in all models tend to capture multiple object parts rather than maintaining a one-to-one correspondence between prototypes and individual semantic parts, which diverges from the intended interpretability objective of prototype-based learning. On average, prototypes focus on approximately one-tenth of the object, as measured by the Object Overlap. This indicates that while a prototype may represent a coherent concept within a single sample image, it typically corresponds to multiple concepts across the test set.
| PIPNet | ProtoPNet | ProtoPNet Pruned | ProtoPool | ||
| (correct pred.) | (correct pred.) | (correct pred.) | (correct pred.) | ||
1. Proto. |
CUB 200 |
2. Proto. |
| (correct pred.) | (correct pred.) | (correct pred.) | (correct pred.) | ||
1. Proto. |
Cars |
2. Proto. |
7.5 Output-Completeness
All output-completeness results in Figure 5 were normalized using either fixed metric boundaries or the maximum values observed across models. As previously mentioned, metric axes are inverted when lower values indicate better performance.
The evaluation of the PRP method for visualizing relevant pixels yields mixed results across all models. The visualizations are most consistent with the feature extraction process of PIPNet, which aligns best with the intended interpretation of high-level object-part extraction, as shown in previous evaluations. This suggests a clear correlation between the accuracy of PRP visualizations and the degree of disentanglement in the extracted features. The second-best visualizations were done on ProtoPNet, although with greater variability in activation location metrics such as PLC and PALC. This indicates that ProtoPNet prototypes are influenced by broader image regions compared to PIPNet, which the PRP method does not fully capture. Examples of PRP visualizations are illustrated in Figure 6 and 7,which show similar sizes of the relevance areas. A similar trend is observed for ProtoPool, where visualizations are less reliable due to pronounced prototype location instability (as seen in the continuity evaluation) and a similar weak inter-prototype contrast, resulting in a more entangled feature space. Overall, these findings suggest that the effectiveness of backpropagation-based visualization methods is strongly dependent on the disentanglement of the learned embedding space.